OpenStudy (anonymous):

What sine function represents an amplitude of 4, a period of pi over 2, no horizontal shift, and a vertical shift of −3?

OpenStudy (campbell_st):

for sin(bx) the period is found by using \[period = \frac{2\pi}{b}\] you have the period and need to find b\[\frac{\pi}{2} = \frac{2\pi}{b}\] solve for b

OpenStudy (campbell_st):

in general, the amplitude is y = asin(x) a = amplitude. You already know the value of a.

OpenStudy (campbell_st):

for the vertical shift its y = asin(bx) + c c is the new centre of the curve. y = sin(x) is centred on y = 0. Moving down 3 it becomes centred on y = -3. Hope it helps